In order to convert a milliammeter of range 1 mA and resistance 1 ohm into a voltmeter of range 10 V, a resistance of how many ohms should be connected?

Grade:12th pass

1 Answers

Khimraj
5 years ago
V = IaRa + IaR
R = V/I– Ra
R = 10/0.001 -1 = 9999 ohm
Hope it clears...................................................